# A Survey on Monocular Re-Localization: From the Perspective of Scene Map Representation

### v1

@@ 6 Vectorized High-Definition Map
$\hat{\textbf{x}}^{\mathcal{G}}_{q}=\mathop{\arg\min}\limits_{\textbf{x}^{\mathcal{G}}_{q}}\sum_{\textbf{m}_{i}\in\mathcal{M}}\sum^{N_{i}}_{j=1}d\left(\textbf{z}^{\mathcal{I}_{q}}_{i(j)},\tilde{\textbf{p}}^{\mathcal{I}_{q}}_{i(j)}\right)$ (19)
where $\textbf{z}^{\mathcal{I}_{q}}_{i(j)}$ is the observation of $j$-th point of $i$-th element $\textbf{m}_{i}$ in image plane $\mathcal{I}_{q}$.
From the basic formulation defined above, HD-MRL methods need to be integrated into a localization system with multiple sensor fusion for acquiring an initial camera pose, which is similar to G-I2P in Sec. 5.1, since HD map elements can be seen as sets of 3D point cloud with semantic information. In this field, Pink et al. Oliver2008CVPRW used aerial images to extract lane markings and their respective positions to build a lane-level map, then matched the lane marking in the image with this pre-built map using ICP algorithm. The subsequent works improve HD-MRL by fusing additional sensors such as GNSS and IMU in an Extended Kalman Filter (EKF) framework. Tao et al. Tao2013IROS proposed to use the EKF fusion framework to integrate lost GPS, dead reckoning, and visual features from lane markings. The lane markings detected by the video camera system provide the lateral distance and heading angle information of a vehicle. Cai et al. Cai2018Sensors also used the matching between 2D visual observation and 3D HD map element as the additional lateral distance observation in the filter-based fusion frameworks.
Figure 14: Examples of alignment between HD map and image semantic segmentation Guo2021IROS . (a) An accurate alignment; (b)-(g) Alignment results under perturbation of freedom: roll, pitch, yaw, x, y and z. Semantic segmentation: freespace (yellow), lane markings (organ and purple). Projection from HD map: lane markings (green) and signboard (cyan).
@@
Another main problem blocking the usage of HD-MRL methods is the the sparsity of HD map landmark features. To solve this problem, multiple sensor fusion often be used in EKF frameworks that IMU and GPS are utilized for continuous localization Tao2013IROS . With the development of VO vins-mono and visual(-inertial) SLAM orbslam ; orbslam2 ; orbslam3 , visual features are shown to be able to serve as reliable constraints for relative pose estimation. They can be considered as complementary localization information in conjunction with the HD map landmark features. Some solutions combine these two pieces of information in a loosely coupled manner, i.e., the HD MRL and the visual feature-based localization results are calculated separately first, and then fused by Kalman Filter Julius2014IV or sliding window-based state estimator MLVHM . However, these loosely-coupled strategies share a common shortcoming: they all require the HD map elements to be minimally self-sufficient for solving vehicle poses, which cannot be satisfied in real-world applications. Wen et al. Wen2020IV ; TM3Loc followed this way to combine VO and HD-MRL but in a tightly-coupled framework with sliding window-based optimization. The basic concept is to optimize the poses with the minimization target consisting of both visual feature residuals from VO and HD map element residuals from HD-MRL, enabling consistent and accurate localization even when HD map elements are insufficient in current observation.
@@
In last two years, some deep learning-based HD-MRL methods have been developed. OrienterNet sarlin2023orienternet estimates 3 DOF pose in a 2D map, OpenStreetMap (OSM), by proposed neural map matching. But its localization accuracy is only limited to sub-meter level due to the low precision of used OSM map. BEV-Locator bevlocator formulates this problem as an E2E learning scheme and proposes a Transformer-based architecture to address the key challenge of the cross-modality matching for HD-MRL. It encodes the map elements from discrete points into structured vectors, and conduct interaction between images and HD map on BEV space. EgoVM he2023egovm extracts BEV features and map embeddings by Transformer Decoder transformer , and then map embeddings are compared with interpolated map features by candidate poses to calculate their similarities, so as to estimate the optimal pose offset. However, BEV-Locator bevlocator and EgoVM he2023egovm needs multiple view images as input, learning-based HD-MRL problem still be unsolved.
@@
Summary: HD map should be the most compact and lightwight scene map, so it is preferred by AD researches. Although it need heavy construction and maintenance cost, the recently proposed crowdsourced map building crowdsourcedmapping and updating GCDL strategies may help to reduce the maintenance cost, making HD map more widely acceptable. MRL in HD map is still a very hard problem because precise data association between visual observations and semantic elements in HD map is theoretically difficult. Traditional methods seek aids from multiple sensor fusion or online odometry while recent learning-based methods leverage advanced representation ability of DNN models. Based on the localization performance and required sensor setup, we believe learning-based method is a more promising way for HD-MRL.
@@
7 Learnt Implicit Map
In the current age of deep learning, scholars begin to rethink the representation format of scene map in a more implicit way. Some recently proposed works encode scene map into neural networks so that the network can achieve amazing things, such as directly recover pose of images (called Absolute Pose Regression, APR), estimate 3D coordinates of each pixel in images (called Scene Coordinate Regression, SCR), or render geometry structure and appearance of scene (called Neural Radiance Field, NeRF). Benefit from high-performance computation device, matured deep learning technology, and tremendous number of training samples, such implicit map may be new choice for MRL methods (called NN-MRL) in the near future. In this section, we review APR, SCR, and NeRF-based MRL methods, as shown in Fig. 15.
7.1 Absolute Pose Regression
@@
APR solution was first proposed by PoseNet posenet to directly regress the position and orientation of the camera with regard to global frame, given a query image $\mathcal{I}_{q}$, by attaching an MLP head to a GoogLeNet backbone googlenet . Although early APR methods achieve less accuracy than the VL-MRL methods (Sec. 4), but enable pose estimation using a single forward pass. The scene map is represented in a totally implicit and very compact formulation, i.e., parameters of DNN $F_{\mbox{APR}}(\cdot)$. Thus, APR method can be simply formatted as:
$\hat{\textbf{x}}^{\mathcal{G}}_{q}=F_{\mbox{APR}}(\mathcal{I}_{q})$ (20)
@@
Single Scene APR (SS-APR): In order to improve the localization accuracy, contemporary APR methods study different designs of CNN backbones Shavit2020CoRR ; Tayyab2017IROS ; hourglassPN , MLP heads Tayyab2017IROS . These models can work very well on training scene but achieve limited accuracy on test scene. Such an overfitting problem is addressed by averaging the predictions of multiple models with randomly dropped activations BayesianPN , or by reducing the dimensional of the global image encoding using Long-Short-Term-Memory (LSTM) layers LSTMPN . Other work focused on the supervision strategy for absolute pose regression to allow adaptive weighting of position- and orientation-associated errors. In original PoseNet posenet , the scale between orientation loss and translation error varies across scenes, e.g., 120 to 750 for indoor scenes while 250 to 2000 for outdoor scenes, which is tricky. Kendall et al. PoseNet2 suggested optimizing the parameters that balance both losses to improve accuracy and avoid manual fine-tuning. This formulation is widely adopted by later APR methods. Alternative orientation representations were also proposed to improve pose loss mapnet . The use of geometric constraints from additional sensors, such as VO and GPS, is also suggested to be formulated as loss terms in network training to improve localization accuracy mapnet . In Saha2018ImprovedVR , SS-APR problem is split into two step that first predicts the most relevant anchor point uniformly defined in the scene and then regress the relative offsets with respect to it. Such a scheme simplifies the APR problem and significantly improve the accuracy. More recently, Wang et al. Wang2019AtLocAG proposed using attention to guide the pose regression process by applying self-attention to the CNN features. CaTiLoc catiloc exploits attention mechanism on full receptive field by Transformer transformer . Xue et al. LSG incorporated learning-based VO estimation and APR in a single DNN and used joint training for RPR and APR so that spatial-temporal consistency can be leveraged to alleviating the uncertainty of pose estimation by augmenting the observation based on the co-visibility from VO estimation. In testing stage, Brahmbhatt et al. mapnet and Xue et al. LSG performed a standard Pose Graph Optimization (PGO) by optimizing only the nodes (APR estimation) with edges (VO estimation) fixed.
@@
Although many modifications are proposed to the architecture and training method originally formulated by Kendall et al. posenet , the main paradigm remained the same that employs (1) a CNN backbone to generate a global latent feature, and (2) a regressor to estimate absolute pose based on extracted features. The backbone and regressor are E2E trained for each scene in the SS-APR methods mentioned above.
@@
Multiple Scene APR (MS-APR): MS-APR methods are later developed aiming to extend a single trained APR model on multiple scenes. Blanton et al. MSPN first proposed Multi-scene PoseNet to first classifies the particular scene related to the query image, and then uses it to index scene-specific regressor to estimate the pose. They train a set of regressors, one per scene, for APR with a set of scene-specific parameterized losses. MS-Transformer MS-Transformer1 ; MS-Transformer2 then learns multi-scene APR with Transformers, where encoders are used to aggregate features with self-attention and decoders transform latent features and scenes encoding into pose predictions. MS-Transformer MS-Transformer1 ; MS-Transformer2 allows the model to focus on informative features for localization by stacked attention, while embedding multiple scenes in parallel.
@@
Summary: Even APR algorithms work well on small scale dataset, e.g., 7Scenes 7scenes , there is no evidence that they can scale to large scale scene. For example, on Cambridge Landmarks dataset posenet that consists of six medium-sized scenes (about 900-5500 $\mbox{m}^{2}$) in an urban environment, the state-of-the-art (SOTA) MS-Transformer MS-Transformer2 achieves an average localization accuracy of 0.98 m and ${}^{\circ}$, much worse than VL-MRL method, InLoc InLoc , with accuracy of 0.11 m and ${}^{\circ}$. APR methods can serve as a initial or coarse pose estimation in some real-time application, and its $O(1)$ computational complexity is also remarkable.
7.2 Scene Coordinate Regression
@@
SCR methods follow the PnP-based solution of MRL methods but in a E2E learning-based way. The scene map $\mathcal{M}$ is also represented as network parameters of DNN $F_{\mbox{SCR}}(\cdot)$. Specifically, SCR methods directly learn to predict 3D scene coordinates $\textbf{p}^{\mathcal{G}}_{i}$ of each 2D pixel $\textbf{p}^{\mathcal{I}_{q}}_{i}$ in the $\mathcal{I}_{q}$ and thus query pose can be solved as a PnP problem based on estimated 2D-3D correspondences, which can be formatted as:
$\begin{split}\{\hat{\textbf{p}}^{\mathcal{G}}_{i}\}&=F_{\mbox{SCR}}(\mathcal{I}_{q})\\ \hat{\textbf{x}}^{\mathcal{G}}_{q}&=\mbox{solvePnP}(\{\textbf{p}^{\mathcal{I}_{q}}_{i},\hat{\textbf{p}}^{\mathcal{G}}_{i}\}_{i})\end{split}$ (21)
@@
where $\{\textbf{p}^{\mathcal{I}_{q}}_{i},\hat{\textbf{p}}^{\mathcal{G}}_{i}\}_{i}$ is a pair of image pixel and estimated scene coordinate.
In this field, early approaches SCRForests ; Abner2014CVPR ; Julien2015CVPR used regression forest models to predict 3D scene coordinates for images, but they need RGB-D images as input. More recent works DSAC ; DSAC2 ; HSCNet ; few-shot-SRC ; DSM ; KFNet ; SANet ; Less_is_More ; Huang2021VSNetVW have replaced regression forests with DNN-based models that only require a RGB image.
Scene-specific SCR (SS-SCR): SCR method is firstly used to estimate dense 3D coordinates for $\mathcal{I}_{q}$ in a previously seen scene, and SCR model need to be specifically trained for each scene, called scene-specific methods. Brachmann et al. DSAC ; DSAC2 ; ESAC ; Less_is_More trained neural networks for this task and combine them with a differentiable RANSAC strategy for MRL. For instance, in ESAC , multiple scene coordinate regression networks are trained as a mixture of experts along with a gating network which assesses the relevance of each expert for a given input, and the final pose estimate is obtained using a novel RANSAC framework, i.e., Expert Sample Consensus (ESAC). Dong et al. few-shot-SRC and Li et al. HSCNet later introduced region classification into their pipelines for effective scene memorization. Huang et al. Huang2021VSNetVW proposed to add a segmentation branch to obtain segmentation on scene-specific landmarks, which can then be associated with 3D landmarks in the scene to estimate camera pose. In Budvytis2019LargeSJ , SS-SCR is formulated as two separate tasks of object instance recognition and local coordinate regression. Since dense 3D structure of scene is required as ground-truth in model training, HSCNet++ Wang2023HSCNetHS removes the dependency on dense ground truth scene coordinates by proposing a pseudo-labelling method, where ground-truth labels at each pixel location are propagated to a fixed spatial neighbourhood. However, SS-SCR need to train models per scene and it commonly pixel-wise ground-truth scene coordinates, the heavy training costs block the wide usage of SS-SCR.
@@
Scene-agnostic SCR (SA-SCR): To boost the generalization ability and scalability of SCR methods, SA-SCR methods are proposed SANet ; DSM ; Revaud2023SACRegSC . They can regress dense scene coordinates based on some given reference views in unseen scenes. SANet SANet first constructs a scene pyramid by extracting multiple scale feature map from reference images, then predicts scene coordinates for the query image with the assistance of scene pyramid. In DSM , authors proposed a Dense Scene Matching (DSM) module. The DSM module predicts scene coordinates by cost volume constructed between query image and reference image with its scene coordinates. Following this way, the Transformer-based SAReg Revaud2023SACRegSC takes a variable number of images (including query images and retrieved reference images) and sparse 2D-3D annotations from SfM database colmap as input.
@@
More implicitly, NeuMap neumap encodes the scene as sparse scene-specific latent codes and estimates scene coordinates with the help of the resulting latent codes. When localizing in a new scene, NeuMap neumap only need to optimize latent code for the new scene without re-training other part of DNN. With the latent code pruning, redundant codes can be removed. So NeuMap neumap can achieves similar performance as DSAC++Less_is_More with a 100 to 1000 times smaller scene map size. Instead of densely estimate scene coordinates, NeuMap neumap and D2S Bui2023D2SRL estimate scene coordinates for sparse 2D key points, making them cost-effectiveness. D2S Bui2023D2SRL also supports updating by self-supervised learning procedure with new observations without requiring camera poses, intrinsic parameters, and ground-truth scene coordinates, which greatly improves the scalability and generalization ability of SCR.
@@
Summary: SCR methods do not explicitly rely on local feature detection, description, and matching, and are able to provide dense correspondences between query image and scene map. They also do not require storing the 3D models and descriptors, as scene map is densely encoded in a learnable DNN. The localization accuracy of SCR is impressively great in small- and medium-scale scenarios 7scenes ; posenet , but its usage in unbounded scene is hard due to expensive training costs.
7.3 Neural Radiance Field

### v3

@@ VI Vectorized High-Definition Map
$\hat{\textbf{x}}^{\mathcal{G}}_{q}=\mathop{\arg\min}\limits_{\textbf{x}^{\mathcal{G}}_{q}}\sum_{\textbf{m}_{i}\in\mathcal{M}}\sum^{N_{i}}_{j=1}d\left(\textbf{z}^{q}_{i(j)},\textbf{p}^{q}_{i(j)}\right)$ (23)
where $\textbf{z}^{q}_{i(j)}$ is the observation of $j$-th point of $i$-th element $\textbf{m}_{i}$ in image plane $\mathcal{I}_{q}$.
From the basic formulation defined above, HD-MRL methods are usually integrated into an online localization system with multiple sensor fusion for acquiring an initial camera pose, which is similar to G-I2P, since HD map elements can be seen as sets of 3D point cloud with semantic information. In this field, Pink et al. [215] built a lane-level map by aerial images, and then matched the lane markings in the query image with the pre-built map using ICP algorithm. However, the sparsity of HD map elements in real-world scenarios restricts the available information for pose estimation, blocking the wide usage of HD-MRL methods. To solve this problem, multiple sensor fusion often be used in Extended Kalman Filter (EKF) frameworks where IMU and GPS are utilized for continuous localization [216, 217]. For the same purpose, recent advances in HD-MRL begin to integrate relative pose constraint by visual feature tracking in VO [218] and visual(-inertial) SLAM [149, 150, 120] so that tracked visual features can be considered as complementary localization information to the HD map elements. In this area, some solutions combine these two pieces of information in a loosely coupled manner. They first perform HD-MRL and VO separately and then fused by Kalman Filter [219] or sliding window-based state estimator [9]. However, such a loosely-coupled strategies require the HD map elements to be minimally self-sufficient for pose estimation, which does not fundamentally address the failure case caused by sparsity of HD map element. So, Wen et al. [220, 8] followed this way to combine VO and HD-MRL but in a tightly-coupled framework with sliding window-based optimization. The optimization target of pose estimation considers both relative constraints of VO and absolute constraint of HD-MRL, enabling consistent and accurate localization even when HD map elements are insufficient in current observation.
Figure 13: Examples of alignment between HD map and image semantic segmentation [221]. (a) An accurate alignment; (b)-(g) Alignment results under perturbation of freedom: roll, pitch, yaw, x, y and z. Semantic segmentation: freespace (yellow), lane markings (organ and purple). Projection from HD map: lane markings (green) and signboard (cyan).
@@
Even with tremendous efforts in traditional HD-MRL methods, the accurate explicit map element matching and pose estimation are still hard problems. In last two years, some DL-based HD-MRL methods have been developed to perform data association in DL feature domains. OrienterNet [230] estimates 3 DoF pose in a 2D map, OpenStreetMap (OSM), by proposed neural map matching. U-BEV [231] leverages a U-Net [232] inspired network to convert surround images and Standard Definition Map (SD Map) into neural representations, i.e., deep CNN features so that 3 DoF pose can be estimated by template matching But their localization accuracy is only limited to sub-meter level due to the low precision of used map. BEV-Locator [233] formulates this problem as an E2E learning scheme and proposes a Transformer-based architecture [65] to address the key challenge of the cross-modality matching for HD-MRL. It encodes the map elements from discrete points into structured vectors, and conduct interaction between images and HD map on BEV space. EgoVM [234] extracts BEV features and map embeddings by Transformer Decoder [65], and then map embeddings are compared with interpolated map features by candidate poses to calculate their similarities in template matching framework, so as to estimate the optimal pose offset. These two methods with HD map achieves centimeter-level localization accuracy, but BEV-Locator [233] and EgoVM [234] needs multiple view images as input, learning-based HD-MRL problem still be unsolved.
@@
Summary: HD map should be the most compact and lightwight scene map, so it is preferred by AD researches. Although it needs heavy construction and maintenance cost, the recently proposed crowdsourced map building [235] and updating [236] strategies may help to reduce the maintenance cost, making HD map more widely acceptable. MRL in HD map is still a very hard problem because precise data association between visual observations and semantic elements in HD map is theoretically difficult. Traditional methods seek aids from multiple sensor fusion or online odometry while recent DL-based methods leverage advanced representation ability of DNN models. Based on the localization performance and required sensor setup, we believe DL-based method is a more promising way for HD-MRL.
VII Learnt Implicit Map
In the current DL age, scholars begin to rethink to represent scene map in a more implicit way. Some recently proposed works encode scene map into neural networks so that the network can achieve amazing things, such as directly recover pose of images (called Absolute Pose Regression, APR), estimate 3D coordinates of each pixel in images (called Scene Coordinate Regression, SCR), or render geometry structure and appearance of scene (called Neural Radiance Field, NeRF), as shown in Fig. 14. Benefit from high-performance computation device, matured DL technology, and sufficient training samples, such implicit map may be a potential choice for MRL methods (called NN-MRL) in the near future.
@@ VII Learnt Implicit Map
VII-A Absolute Pose Regression
@@
APR solution was first proposed by PoseNet [27] to directly regress the 6 DoF pose $\textbf{x}^{\mathcal{G}}_{v}$ of the camera with regard to global frame $\mathcal{G}$, given a query image $\mathcal{I}_{q}$. The original PoseNet [27] is composed by a feature encoder that is a GoogLeNet backbone [237], and a pose regressor that is a MLP head. Although early APR methods achieve less accuracy than the VL-MRL methods, they enable pose estimation using a single forward pass and remove the need of saving scene map. The scene map is represented in a totally implicit and very compact formulation, i.e., parameters of DNN $\mbox{APR}(\cdot)$. Thus, APR method can be simply formatted as:
$\hat{\textbf{x}}^{\mathcal{G}}_{q}=\mbox{APR}(\mathcal{I}_{q})$ (24)
@@
Single Scene APR (SS-APR): In order to improve the localization accuracy, contemporary APR methods achieve further improvements by developing different CNN backbones [238, 239, 240] or MLP heads [239], modelling localization uncertainty [241], or introducing Long-Short-Term-Memory (LSTM) layers [242] in the networks. Since the scale of orientation (in degrees) and position (in meters) are quite different across scenes, APR methods need to use weight to balance these two factors. In original PoseNet [27], the scale between orientation error and translation error varies in different scenes, e.g., 120 to 750 for indoor scenes while 250 to 2000 for outdoor scenes, which is tricky. Kendall et al. [243] proposed a new loss function with learnable weight of orientation error and translation error so that manual fine-tunning can be avoided, which is popularly used in later works. In [244], authors improved by using logarithm of a unit quaternion instead of traditional unit quaternion as a more advanced orientation representation, and introduced geometric constraints from additional sensors, such as VO and GPS, as new loss terms in network training to improve the performance. In [245], SS-APR problem is split into two step that first predicts the most relevant anchor point uniformly defined in the scene and then regress the relative offsets with respect to it. Such a scheme simplifies the APR problem and significantly improve the accuracy. For higher localization accuracy, Wang et al. [246] applied self-attention to the CNN features to guide the pose regression while CaTiLoc [247] exploits attention mechanism on full receptive field by Transformer [65]. Xue et al. [248] incorporated L-RPR and APR in a single DNN and jointly trained the network so that the uncertainty of APR can be alleviated by augmenting the observation based on the co-visibility from VO estimation. In testing stage, Brahmbhatt et al. [244] and Xue et al. [248] performed a standard Pose Graph Optimization (PGO) by optimizing only the nodes (APR estimation) with edges (VO or RPR estimation) fixed.
@@
Although many modifications are proposed to the architecture and training method originally formulated by PoseNet [27], the main paradigm remained the same that employs 1) a CNN backbone to generate features, which is considered to contain pose-related information, and 2) a regressor to estimate absolute pose based on extracted features. The backbone and regressor are E2E trained for each scene in the SS-APR methods mentioned above.
@@
Multiple Scene APR (MS-APR): MS-APR methods are later developed aiming to extend a single trained APR model on multiple scenes. Blanton et al. [249] first proposed multi-scene PoseNet to first classifies the scene where the query image is taken, and then retrieve a scene-specific regressor corresponding to this scene for pose regression. They train a set of regressors, one per scene, for APR with a set of scene-specific parameterized losses. MS-Transformer [250, 251] then learn multi-scene APR with a single unified Transformers-based network [65], where scene-specific queries (scene encodings) are used in Transformer [65], so they can encode many scenes in parallel and achieve MS-APR. During pose regression, both CNN features and scenes encodings are considered.
@@
Summary: Although APR algorithms have been developed for almost 10 years, there is no evidence that they can scale to large scale scene. For example, on Cambridge Landmarks dataset [27] that consists of six medium-sized scenes (about 900-5500 $\mbox{m}^{2}$) in an urban environment, the SOTA MS-Transformer [251] only achieves an average localization accuracy of 0.98 m and ${}^{\circ}$. APR methods can serve as an initial or coarse pose estimation in some real-time application, and its $O(1)$ computational complexity is also remarkable.
VII-B Scene Coordinate Regression
@@
SCR methods follow the PnP-based solution of MRL methods but in a E2E learning-based way. The scene map $\mathcal{M}$ is also represented as network parameters of DNN $\mbox{SCR}(\cdot)$. Specifically, SCR methods directly learn to predict 3D scene coordinates $\textbf{P}_{i}$ of each 2D pixel $\textbf{p}^{q}_{i}$ in the $\mathcal{I}_{q}$ and thus query pose can be solved as a PnP problem based on estimated 2D-3D correspondences:
$\begin{split}\{\hat{\textbf{P}}_{i}\}&=\mbox{SCR}(\mathcal{I}_{q})\\ \hat{\textbf{x}}^{\mathcal{G}}_{q}&=\mbox{SolvePnP}(\{\textbf{p}^{q}_{i},\hat{\textbf{P}}_{i}\}_{i})\end{split}$ (25)
@@
where $\{\textbf{p}^{q}_{i},\hat{\textbf{P}}_{i}\}_{i}$ is a pair of image pixel and estimated scene coordinate.
In this field, early approaches [252, 253, 254] used regression forest models and they required RGB-D images as input. More recent works [255, 256, 257, 258, 259, 260, 43, 261, 262] have instead applied DNN-based models fed by RGB images, which greatly improves the usability of the SCR solution.
Scene-specific SCR (SS-SCR): SCR method is firstly used to estimate dense 3D coordinates for $\mathcal{I}_{q}$ in a previously seen scene, and it need to be specifically trained for each scene, called scene-specific methods. To reduce the localization uncertainty caused by imperfect network estimation, the PnP solver is always integrated in a RANSAC framework so that an estimation with the most inliers will be selected. However, vanilla RANSAC is non-differentiable, blocking E2E training for SCR models. So, Brachmann et al. [255, 256, 263, 261] proposed differentiable RANSAC strategies for SCR so that pose can be E2E estimated. For instance, in DSAC [255, 256], Brachmann et al. scored estimated scene coordinates (can be seen as 2D-3D correspondences) using a CNN and selected final pose by probabilistic selection while in [263], authors proposed a novel RANSAC framework called Expert Sample Consensus (ESAC) that many SCR networks are regard as a set of “experts” and a gating network is trained to measure the relevance of each expert regarding $\mathcal{I}_{q}$.
For effectively encode scene map in SCR model, some works turn to learn region-wise label for each pixel. HSCNet [257] predicts scene coordinates in a coarse-to-fine manner that it first classifies pixels into corresponding sub-regions and predict scene coordinates to reduce prediction ambiguity. Such a joint classification and regression strategy is then well studied [264, 258, 262, 265]. This way is concluded to alleviate the dependency on dense ground-truth scene coordinates in SCR model training [265]. Specifically, HSCNet++ [265] proposes a pseudo-labelling method, where ground-truth labels at each pixel location are propagated to a fixed spatial neighbourhood. However, SS-SCR trains models per scene and commonly requires pixel-wise ground-truth scene coordinates. The heavy training costs of SS-SCR block its developments.
@@
Scene-agnostic SCR (SA-SCR): To boost the generalization ability and scalability of SCR methods, SA-SCR methods are proposed [43, 259, 266]. They can regress dense scene coordinates based on some given reference views in unseen scenes. SANet [43] constructs a scene pyramid by extracting MS feature map from reference images, then predicts scene coordinates for the query image with the assistance of the scene pyramid. In [259], authors proposed a Dense Scene Matching (DSM) module. The DSM module predicts scene coordinates by cost volume constructed between query image and reference image with its scene coordinates. Following this way, the Transformer-based SAReg [266] takes a variable number of images (including query images and retrieved reference images) and sparse 2D-3D annotations from SfM database [119] as inputs.
@@
More implicitly, NeuMap [28] encodes the scene as sparse scene-specific latent codes and estimates scene coordinates with the help of the resulting latent codes. When localizing in a new scene, NeuMap [28] only need to generate new latent code for the new scene without re-training other parts of network. By pruning redundant latent code, NeuMap [28] can achieves similar performance as DSAC++[261] with a 100 to 1000 times smaller scene map size. Instead of densely estimate scene coordinates, NeuMap [28] and D2S [267] estimate scene coordinates for sparse 2D key points in query images, making them effective. D2S [267] also supports updating by self-supervised learning procedure with new observations without requiring camera poses, intrinsic parameters, and ground-truth scene coordinates, which greatly improves the scalability and generalization ability of SCR.
@@
Summary: SCR methods do not explicitly rely on local feature extraction and matching, and they are able to provide dense correspondences between query image and scene map. They also do not require storing the 3D models and descriptors, as scene map is implicitly encoded in a learnable DNN. The localization accuracy of SCR is impressively great in small- and medium-scale scenarios [268, 27], but the expensive training costs block its usage in unbounded scene.
VII-C Neural Radiance Field